OpenStudy (anonymous):

How do i work out 4x^2 - 9 ? Thanks!

OpenStudy (anonymous):

what are you trying to do with it? Or what are you finding?

OpenStudy (anonymous):

I need to factorise that expression to find values of X

OpenStudy (anonymous):

oh... difference of squares... so... (2x+3)(2x-3)... then you set each = to 0.

OpenStudy (anonymous):

so 2x+3=0 so 2x=-3 so x=-3/2

OpenStudy (anonymous):

Did you just do (2x+3)^2 = 4x^2-9? How did you get the (2x-3)

OpenStudy (anonymous):

Thanks!!!!

OpenStudy (anonymous):

difference of squares. if you foil out (2x+3)^2 you get 4x^2+12x+9

OpenStudy (anonymous):

youre welcome :D

OpenStudy (anonymous):

Ps, how do i end this post/ say that the question has been solved, i dont wanna waster anyones time!

OpenStudy (anonymous):

lol umm... look at the place in the left column of the page where u asked the question and press the close question button :)